OpenStudy (anonymous):

how do you graph f(x)=tan4x

OpenStudy (anonymous):

Hey so the graph of \(f(x)=tan(4x)\) Is, of course, related to the graph of \(f(x)=tan(x)\). The difference is that there exists a co-efficient of four in front of the x. This means that all the graph of \(tan(x)\) will be compressed horizontally by a factor of \(\frac{1}{4}\) So it will be the same with the exception that instead of repeating every \(\pi\), it will repeat every \(\frac{\pi}{4}\)